OpenStudy (anonymous):

A system of equations is shown below: x + 3y = 5 (equation 1) 7x - 8y = 6 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

Show that the solution to the system of equations 3x + y = 5 and 8x -7y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations 8x - 5y = 11 and 7x - 8y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations 15x + 13y = 17 and 7x - 8y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations -13x + 15y = 17 and 7x - 8y = 6 is the same as the solution to the given system of equations

OpenStudy (anonymous):

@TwiztTiez

OpenStudy (twizttiez):

Okaii

OpenStudy (twizttiez):

Well lets get started

OpenStudy (anonymous):

:3

OpenStudy (twizttiez):

I think your answer would be D the bottom one Each answer choice has two equations: The second choice is the same for all and we can ignore it. The first choice is: "equation 1 is replaced with the sum of equation 1 and a multiple of equation 2"

OpenStudy (twizttiez):

So taking that we know that From the first choice in each answer, subtract equation 1 and see if what is left is a multiple of equation 2. If yes then that is the answer.

OpenStudy (twizttiez):

Make sense?